Solve y+px=p²x⁴


1
Expert's answer
2021-08-25T18:26:19-0400

The given differential equation is


"y+px=p^2x^4"

Differentiating with respect to "x," we get 


"\dfrac{dy}{dx}+p+x\dfrac{dp}{dx}=2px^4\dfrac{dp}{dx}+4p^2x^3"

"p+p+x\dfrac{dp}{dx}=2px^4\dfrac{dp}{dx}+4p^2x^3"

"\dfrac{dp}{dx}(x-2px^4)=4p^2x^3-2p"

"x(1-2px^3)\dfrac{dp}{dx}=-2p(1-2px^3)"

"1-2px^3=0"

"p=\dfrac{1}{2x^3}"

Substitute


"y+(\dfrac{1}{2x^3})x=(\dfrac{1}{2x^3})^2x^4"

"y=-\dfrac{1}{2x^2}+\dfrac{1}{4x^2}"

"y=-\dfrac{1}{4x^2}"

"x\dfrac{dp}{dx}=-2p"

"\dfrac{dp}{p}+2\dfrac{dx}{x}=0"

Integrating, we get


"\ln|p|+\ln x^2=\ln C"

"px^2=C"

Substitute


"px=\dfrac{C}{x}"


"(px^2)^2=p^2x^4=C^2"


"y+px=p^2x^4"

Then


"y+\dfrac{C}{x}=C^2"